# Towards Unified Latent Space for 3D Molecular Latent Diffusion Modeling

### v2

@@ Abstract
3D molecule generation is crucial for drug discovery and material science, requiring models to process complex multi-modalities, including atom types, chemical bonds, and 3D coordinates. A key challenge is integrating these modalities of different shapes while maintaining SE(3) equivariance for 3D coordinates. To achieve this, existing approaches typically maintain separate latent spaces for invariant and equivariant modalities, reducing efficiency in both training and sampling. In this work, we propose Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling (UAE-3D), a multi-modal VAE that compresses 3D molecules into latent sequences from a unified latent space, while maintaining near-zero reconstruction error. This unified latent space eliminates the complexities of handling multi-modality and equivariance when performing latent diffusion modeling. We demonstrate this by employing the Diffusion Transformer–a general-purpose diffusion model without any molecular inductive bias–for latent generation. Extensive experiments on GEOM-Drugs and QM9 datasets demonstrate that our method significantly establishes new benchmarks in both de novo and conditional 3D molecule generation, achieving leading efficiency and quality.
Towards Unified Latent Space for 3D Molecular Latent Diffusion Modeling
Yanchen Luo1, Zhiyuan Liu2, Yi Zhao1, Sihang Li1, Kenji Kawaguchi2, Tat-Seng Chua2, Xiang Wang1 1University of Science and Technology of China 2National University of Singapore luoyanchen@mail.ustc.edu.cn
@@ 1 Introduction
The discovery of novel molecules is fundamental to advancements across various scientific fields, particularly in drug and material development. Given significant progress has been made in designing 2D molecular graphs Jin et al. (2018); Irwin et al. (2022); Polykovskiy et al. (2020); Shi et al. (2020), recent research has increasingly focused on the generation of 3D molecules Hoogeboom et al. (2022); Huang et al. (2023b). Unlike 2D molecular generation, which focuses on forming valid molecular structures based on chemical bonds, 3D generation must also predict 3D atomic coordinates that align with the 2D structure. Accurate 3D molecule generation is essential to power many important applications, such as structure-based drug design Zhang et al. (2023) and inverse molecular design targeting quantum chemical properties Bao et al. (2023).
Figure 1: (a) A 3D molecule is characterized by features of distinct modalities. (b) Conventional approaches use separate latent spaces for equivariant (3D coordinates) and invariant (2D features) components, leading to fragmented molecular representations. (c) Our UAE-3D establishes a unified SE(3)-equivariant latent space that jointly encodes both modalities
Table 1: Comparing the reconstruction error between UAE-3D and GeoLDM’s VAE. Atom/Bond accuracy measures correct atom/bond type predictions. Coordinate RMSD quantifies geometric deviation from the ground truth 3D coordinates.
Metric GeoLDM Xu et al. (2023) UAE-3D (Ours)
Atom Accuracy (%) 98.6 100.0
Bond Accuracy (%) 96.2 100.0
Coordinate RMSD (Å) 0.1830 0.0002
@@
3D molecule generation is challenging due to its multi-modal nature. As shown in Figure 1(a), a 3D molecule consists of features of three distinct modalities: atom types, atomic coordinates, and chemical bonds. This requires the generation model to handle both discrete (e.g., atom types) and continuous features (e.g., coordinates), while also addressing differences in feature shapes (atom-wise v.s. edge-wise features). Worse still, the modality of 3D coordinate requires special care to respect rotational and translational equivariance, whereas other modalities do not, further complicating the task. Mitigating this multi-modal challenge, prior works mostly process each modality in separate latent spaces: some maintain separate diffusion processes for each modality Huang et al. (2023b); Vignac et al. (2023); Le et al. (2024); Hua et al. (2023), while others interleave the prediction of different modalities across autoregressive generation steps Gebauer et al. (2019, 2021). However, we argue that handling each modality separately complicates the model design, reducing both the training and sampling efficiency. Moreover, processing each modality separately risks compromising the consistency between them. These issues raise a critical question: can we design a unified generative model that seamlessly integrates all three modalities of 3D molecule generation?
@@
To answer this research question, we propose to build a multi-modal latent diffusion model (LDM) Rombach et al. (2022) for the unified generative modeling of 3D molecules. LDM extends the diffusion paradigm Ho et al. (2020) by operating in a compressed latent space learned through variational autoencoders (VAEs) Kingma and Welling (2014), offering improved computational efficiency and generation quality. For 3D molecules, we can build a multi-modal VAE that compresses all three modalities in a single unified latent space. Scrutinizing the previous works, we identify the following challenges to achieve this purpose:
@@
  • Building a Single Unified Latent Space for 3D Molecules. While previous studies have explored LDMs for 3D molecule and protein generation Xu et al. (2023); Kong et al. (2024), they fail to build a unified latent space that integrates all modalities. Their difficulty arises from the reliance on neural networks with baked-in 3D equivariance Satorras et al. (2021); Huang et al. (2023b), which mostly maintain separate latent spaces for 3D equivariant and invariant modalities (cf. Figure 1(b)).
@@
  • Ensuring Near-Lossless Compression. Maintaining a low reconstruction error for the multi-modal VAE is critical, as even small errors can result in invalid or unstable 3D structures. Furthermore, imprecise latents that cannot reproduce the original molecule can propagate their errors to the LDM, easily disrupting the generative modeling. Despite its critical importance, reconstruction error has been largely overlooked in previous works Xu et al. (2023); Kong et al. (2024) (cf. Table 1).
@@
In this work, we introduce Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling (UAE-3D), a VAE that can compress the multi-modal features of 3D molecules into a unified latent space while maintaining near zero (100% atom/bond accuracy and 2E-4 coordinate RMSD) reconstruction error. To obtain a latent space integrating both 3D equivariant and invariant modalities, we draw inspiration from the “bitter lesson” Wang et al. (2024) of 3D molecules: rather than using models with intricate, baked-in 3D equivariance, UAE-3D trains a neural network to “learn” 3D equivariance through our tailor-made SE(3) augmentations, encouraging the transformation on the input coordinates are reflected equivariantly on the output coordinates. Moreover, UAE-3D employs the Relational Transformer Diao and Loynd (2023) as its encoder, leveraging its scalability and flexibility to incorporate both atom-wise and edge-wise features. A Transformer-based Vaswani et al. (2017) decoder is jointly trained with the encoder to reconstruct both 3D invariant and equivariant molecular features.
@@
Despite its simplicity, UAE-3D can compress 3D molecules as token sequences in a unified latent space, eliminating the complexities of handling multi-modalities and 3D equivariance in latent diffusion modeling. As shown in Table 1, UAE-3D reaches lossless reconstruction of atom and bond types, with a near-zero coordinate RMSD. To further demonstrate its effectiveness, we employ the Diffusion Transformer (DiT) Peebles and Xie (2023), a general-purpose diffusion backbone without any molecular inductive bias, to model UAE-3D’s latents. We show that DiT can successfully generate stable and valid 3D molecules, with significantly improved training (by 2.7 times) and sampling efficiency (by 7.3 times). We refer to this LDM pipeline as UDM-3D: Unified Latent Diffusion Modeling for 3D Molecule Generation.
Extensive experiments demonstrate that UDM-3D, powered by UAE-3D, achieves state-of-the-art results in both de novo and conditional 3D molecule generation on the QM9 Ramakrishnan et al. (2014) and GEOM-Drugs Axelrod and Gómez-Bombarelli (2020) datasets. Ablation studies show the effectiveness of our key components.
Figure 2: Overview of the UDM-3D and UAE-3D models. The UAE-3D encodes 3D molecules from molecular space into a unified latent space, integrating multi-modal features such as atom types, chemical bonds, and 3D coordinates. This unified latent space is then utilized by the UDM-3D, which employs a DiT to perform generative modeling. Then the denoised latents are decoded back into 3D molecules.
@@ 2 Background: 3D Molecular Latent Diffusion Models
In this work, we explore latent diffusion models (LDMs) for 3D molecule generation. A 3D molecular LDM involves a 3D molecular VAE to compress 3D molecules into the latent space, where a diffusion model performs generative modeling. Below, we briefly introduce both components.
2.1 Molecular Variational Auto-Encoding
@@
A molecular VAE consists of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. Given a 3D molecule $\mathbf{G}$, the encoder $\mathcal{E}$ maps it into a sequence of latent tokens $\mathcal{E}(\mathbf{G})=\mathbf{Z}=\{\boldsymbol{z}_{i}\in\mathbb{R}^{d}|i\in\mathcal{V}\}$. Each latent $\boldsymbol{z}_{i}$ is sampled from a Gaussian distribution $\mathcal{N}(\boldsymbol{z}_{i};\boldsymbol{\mu}_{i},\boldsymbol{\sigma}_{i})$ using the reparameterization trick Kingma and Welling (2014), where $\boldsymbol{\mu}_{i}$ and $\boldsymbol{\sigma}_{i}$ are learned parameters for atom $i$. The decoder $\mathcal{D}$ then reconstructs $\mathbf{G}$ from these latents: $\hat{\mathbf{G}}=\mathcal{D}(\mathbf{Z})$. The VAE is trained to minimize a reconstruction loss and a regularization term to ensure the latent space follows a standard Gaussian prior:
$\mathcal{L}_{\operatorname{VAE}}=$ (1)
$\mathbb{E}_{\mathbf{G}\sim p_{\operatorname{data}}}\left[\|\mathbf{G}-\hat{\mathbf{G}}\|+D_{\operatorname{KL}}(q(\mathbf{Z}|\mathbf{G})\|p(\boldsymbol{z}))\right],$
where $D_{\operatorname{KL}}$ denotes the Kullback-Leibler divergence; $q(\mathbf{Z}|\mathbf{G})=\prod_{i\in\mathcal{V}}\mathcal{N}(\mathbf{z}_{i};\boldsymbol{\mu}_{i},\boldsymbol{\sigma}_{i})$ is the approximated posterior distribution; and $\|\mathbf{G}-\hat{\mathbf{G}}\|$ measures the difference between the original and the predicted graph, whose definition varies across different works. In Xu et al. (2023), $\|\mathbf{G}-\hat{\mathbf{G}}\|$ combines an MSE loss for 3D coordinates and a cross-entropy loss for atom types.
2.2 Diffusion Model
Building on the VAE’s latent $\mathbf{Z}$ for 3D molecules, a molecular LDM performs generative modeling with a diffusion model Rombach et al. (2022); Kingma et al. (2021). In the forward diffusion process, we gradually add noise to the original latent $\mathbf{Z}^{(0)}=\mathbf{Z}$ following:
$\mathbf{Z}^{(t)}\sim\mathcal{N}(\mathbf{Z}^{(t)};\sqrt{\bar{\alpha}^{(t)}}\mathbf{Z}^{(0)},(1-\bar{\alpha}^{(t)})\mathbf{I}),$ (2)
where $t\in(0,1]$ is the diffusion timestep, and $\bar{\alpha}^{(t)}$ is a hyperparameter controlling the signal and noise ratio. At the final step ($t=1$), $\mathbf{Z}^{(1)}$ approaches the prior distribution, an isotropic Gaussian distribution: $\mathbf{Z}^{(1)}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})$. In practice, $\mathbf{Z}^{(t)}$ is sampled as $\mathbf{Z}^{(t)}=\sqrt{\bar{\alpha}^{(t)}}\mathbf{Z}^{(0)}+\sqrt{1-\bar{\alpha}^{(t)}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})$.
Given the noised latents $\mathbf{Z}^{(t)}$, a diffusion model $\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},t)$ trained to predict the added noise $\boldsymbol{\epsilon}$ by minimizing the MSE loss $\|\epsilon_{\theta}(\mathbf{Z}^{(t)},t)-\boldsymbol{\epsilon}\|^{2}$. Once trained, new latents $\hat{\mathbf{Z}}^{(0)}$ are sampled using the diffusion model $\boldsymbol{\epsilon}_{\theta}$ via the iterative denoising process Ho et al. (2020), and the decoder $\mathcal{D}$ reconstructs the corresponding 3D molecule via $\hat{\mathbf{G}}=\mathcal{D}(\hat{\mathbf{Z}}^{(0)})$.
@@ 3 Methodology
3.1 UAE-3D: Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling
UAE-3D is designed to address the complexities of multi-modal and equivariance of molecular data by compressing the atomic features, bond features, and atomic coordinates into a unified latent space. This is achieved through three key components: (1) a Relatinoal-Transformer Diao and Loynd (2023) that effectively integrates the multi-modal features into token sequences, (2) a Transformer-based Vaswani et al. (2017) decoder and the reconstruction loss for 3D molecule reconstruction; and (3) our tailor-made data augmentations to train the model to learn 3D equivariance.
Compressing 3D Molecules with Relational Transformer (R-Trans).
Given a 3D molecular $\mathbf{G}=\langle\mathbf{F},\mathbf{E},\mathbf{X}\rangle$, we compute its initial atom-wise and edge-wise embeddings as:
$\mathbf{H}^{n}$ $=\operatorname{MLP}([\mathbf{X};\mathbf{F}])$ $\in\mathbb{R}^{|\mathcal{V}|\times d},$ (3)
$\mathbf{H}^{e}$ $=\operatorname{MLP}([\mathbf{E};\mathbf{D}])$ $\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times d},$ (4)
$\mathbf{D}_{ij}$ $=\operatorname{GBF}(\|\mathbf{X}_{i}-\mathbf{X}_{j}\|^{2})$ $\in\mathbb{R}^{d},$ (5)
where $\operatorname{GBF}(\cdot)$ implements the Gaussian basis functions to expand inter-atomic distances as feature vectors Thölke and Fabritiis (2022); $d$ is the embedding dimension. The initial node embeddings $\mathbf{H}^{n}$ combines both atomic features and its 3D coordinates, while the edge embeddings $\mathbf{H}^{e}$ incorporate bond features and inter-atomic distances. $\mathbf{H}^{n}$ and $\mathbf{H}^{e}$ together effectively represent all 3D molecular information for the subsequent encoding. After obtaining the initial embeddings, we process them through $L$ layers of R-Trans blocks Diao and Loynd (2023):

### v3

# Towards Unified and Lossless Latent Space for 3D Molecular Latent Diffusion Modeling

@@ Abstract
3D molecule generation is crucial for drug discovery and material science, requiring models to process complex multi-modalities, including atom types, chemical bonds, and 3D coordinates. A key challenge is integrating these modalities of different shapes while maintaining SE(3) equivariance for 3D coordinates. To achieve this, existing approaches typically maintain separate latent spaces for invariant and equivariant modalities, reducing efficiency in both training and sampling. In this work, we propose Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling (UAE-3D), a multi-modal VAE that compresses 3D molecules into latent sequences from a unified latent space, while maintaining near-zero reconstruction error. This unified latent space eliminates the complexities of handling multi-modality and equivariance when performing latent diffusion modeling. We demonstrate this by employing the Diffusion Transformer–a general-purpose diffusion model without any molecular inductive bias–for latent generation. Extensive experiments on GEOM-Drugs and QM9 datasets demonstrate that our method significantly establishes new benchmarks in both de novo and conditional 3D molecule generation, achieving leading efficiency and quality. On GEOM-Drugs, it reduces FCD by 72.6% over the previous best result, while achieving over 70% relative average improvements in geometric fidelity.
11footnotetext: Corresponding authors.
@@ 1 Introduction
The discovery of novel molecules is fundamental to various scientific fields, particularly in drug and material development. Given significant progress has been made in designing 2D molecular graphs [1, 2, 3, 4], recent research has increasingly focused on the generation of 3D molecules [5, 6]. Unlike 2D molecular generation, which focuses on forming valid molecular structures based on chemical bonds, 3D generation must also predict 3D atomic coordinates that align with the 2D structure. Accurate 3D molecule generation is essential to power many important applications, such as structure-based drug design [7] and inverse molecule design targeting quantum properties [8].
Figure 1: (a) A 3D molecule has multi-modal features. (b) Prior methods use separate latent spaces for equivariant (3D) and invariant (2D) modalities, inducing unnecessary complexity for the model architecture. (c) UAE-3D reduces this complexity by establishing a unified and lossless latent space that integrates all molecular modalities.
Figure 2: Comparing UAE-3D and UDM-3D with other methods on the QM9 dataset. (a;b) Reconstruction errors on the test set. (c;d) Comparing training and inference time.
@@
3D molecule generation is challenging due to its multi-modal nature. As shown in Figure 2(a), a 3D molecule consists of features of three distinct modalities: atom types, atomic coordinates, and chemical bonds. This requires the generation model to handle both discrete (e.g., atom types) and continuous features (e.g., coordinates), while also addressing differences in feature shapes (atom-wise v.s. edge-wise features). Worse still, the modality of 3D coordinate requires special care to respect rotational and translational equivariance, whereas other modalities do not, further complicating the task. Mitigating this multi-modal challenge, prior works mostly process each modality in separate latent spaces: some maintain separate diffusion processes for each modality [6, 9, 10, 11], while others interleave the prediction of different modalities across autoregressive generation steps [12, 13]. However, we argue that handling each modality separately complicates the model design, reducing both the training and sampling efficiency. Moreover, processing each modality separately risks compromising the consistency between them. These issues raise a critical question: Can we design a unified generative model that seamlessly integrates all three modalities of 3D molecule generation?
@@
To answer this research question, we propose to build a multi-modal latent diffusion model (LDM) [14] for the unified generative modeling of 3D molecules. LDM extends the diffusion paradigm [15] by operating in a compressed latent space learned through variational autoencoders (VAEs) [16], offering improved computational efficiency and generation quality. For 3D molecules, we can build a multi-modal VAE that compresses all three modalities in a single unified latent space. Scrutinizing the previous works, we identify the following challenges to achieve this purpose:
@@
  • The Challenge of Compressing Multi-modality in One Latent Space. While previous studies have explored LDMs for 3D molecule and protein generation [17, 18], they fail to build a unified latent space that integrates all modalities. Their difficulty arises from the reliance on neural networks with baked-in 3D equivariance [19, 6], which mostly maintain separate latent spaces for 3D equivariant and invariant modalities (cf. Figure 2(b)).
@@
  • The Challenge of Ensuring Near-Lossless Compression. Ensuring a low reconstruction error for the molecular VAE is critical, as even small errors can result in invalid or unstable 3D structures. Furthermore, imprecise latents that cannot reproduce the original molecule can propagate their errors to the LDM, easily disrupting the generative modeling. Despite its critical importance, reconstruction error has been largely overlooked in previous works [17, 18] (cf. Figure 2(a;b)).
@@
In this work, we introduce UAE-3D, Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling, a VAE that can compress the multi-modal features of 3D molecules into a unified latent space while maintaining near zero (100% atom/bond accuracy and 2E-4 coordinate RMSD) reconstruction error. To obtain a latent space integrating both 3D equivariant and invariant modalities, we draw inspiration from the “bitter lesson” [20] of 3D molecules: rather than using models with intricate, baked-in 3D equivariance, UAE-3D trains a neural network to “learn” 3D equivariance through our tailor-made SE(3) augmentations, encouraging the transformation on the input coordinates to be reflected equivariantly on the output coordinates. Moreover, UAE-3D employs the Relational Transformer [21] as its encoder, leveraging its scalability and flexibility to incorporate both atom-wise and edge-wise features. A Transformer-based [22] decoder is jointly trained with the encoder to reconstruct both 3D invariant and equivariant molecular features.
@@
Despite its simplicity, UAE-3D can compress 3D molecules into token sequences in a unified latent space, eliminating the complexities of handling multi-modalities and 3D equivariance in latent diffusion modeling. Figure 2(a;b) show that UAE-3D reaches lossless reconstruction of atom and bond types, with a near-zero coordinate RMSD. To further demonstrate its effectiveness, we employ the Diffusion Transformer (DiT) [23], a general-purpose diffusion backbone without any molecular inductive bias, to model UAE-3D’s latents. We show that DiT can successfully generate stable and valid 3D molecules, with significantly improved training (by 2.7 times) and sampling (by 7.3 times) efficiency (cf. Figure 2(c;d)). We refer to this LDM pipeline as UDM-3D: Unified Latent Diffusion Modeling for 3D Molecule Generation.
Extensive experiments demonstrate that UDM-3D, powered by UAE-3D, achieves state-of-the-art results in both de novo and conditional 3D molecule generation on the QM9 [24] and GEOM-Drugs [25] datasets. On GEOM-Drugs, our model achieves a 72.6% reduction in Fréchet ChemNet Distance [3], while significantly reducing the MMD of bond length, bond angle, and dihedral angle by 88.4%, 55.6%, and 74.0%, respectively. In conditional generation on QM9, it achieves the lowest MAE in five out of six properties. These results confirm the effectiveness of our unified modeling approach. Ablation studies show the effectiveness of our key components. Further analysis reveals that UAE-3D’s latents present structured variations across geometric movement.
@@ 2 Background: 3D Molecular Latent Diffusion Models
A 3D molecular LDM involves a 3D molecular VAE to compress 3D molecules into the latent space, where a diffusion model performs generative modeling. Below, we introduce both components.
@@
3D Molecular Variational Auto-Encoding. A molecular VAE consists of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. Given a 3D molecule $\mathbf{G}$, the encoder $\mathcal{E}$ maps it into a sequence of latent tokens $\mathcal{E}(\mathbf{G})=\mathbf{Z}=\{\boldsymbol{z}_{i}\in\mathbb{R}^{d}|i\in\mathcal{V}\}$. Each latent $\boldsymbol{z}_{i}$ is sampled from a Gaussian distribution $\mathcal{N}(\boldsymbol{z}_{i};\boldsymbol{\mu}_{i},\boldsymbol{\sigma}_{i})$ using the reparameterization trick [16], where $\boldsymbol{\mu}_{i}$ and $\boldsymbol{\sigma}_{i}$ are learned parameters for atom $i$. The decoder $\mathcal{D}$ then reconstructs $\mathbf{G}$ from these latents: $\hat{\mathbf{G}}=\mathcal{D}(\mathbf{Z})$. The VAE is trained with a reconstruction loss and a regularization term:
$\mathcal{L}_{\operatorname{VAE}}=\mathbb{E}_{\mathbf{G}\sim p_{\operatorname{data}}}\left[\|\hat{\mathbf{G}}-\mathbf{G}\|+D_{\operatorname{KL}}(q(\mathbf{Z}|\mathbf{G})\|p(\boldsymbol{z}))\right],$ (1)
where $D_{\operatorname{KL}}$ denotes the Kullback-Leibler divergence; $q(\mathbf{Z}|\mathbf{G})=\prod_{i\in\mathcal{V}}\mathcal{N}(\mathbf{z}_{i};\boldsymbol{\mu}_{i},\boldsymbol{\sigma}_{i})$ is the approximated posterior distribution; and $\|\hat{\mathbf{G}}-\mathbf{G}\|$ measures the difference between the original and the predicted graph, whose definition varies across different works. In [17], $\|\mathbf{G}-\hat{\mathbf{G}}\|$ combines an MSE loss for 3D coordinates and a cross-entropy loss for atom types.
Diffusion Model. Building on the molecular VAE’s latent $\mathbf{Z}$, a diffusion model [14, 26] performs generative modeling. In forward diffusion process, we gradually add noise to the original latent $\mathbf{Z}^{(0)}=\mathbf{Z}$ following $\mathbf{Z}^{(t)}\sim\mathcal{N}(\mathbf{Z}^{(t)};\sqrt{\bar{\alpha}^{(t)}}\mathbf{Z}^{(0)},(1-\bar{\alpha}^{(t)})\mathbf{I}),$ where $t\in(0,1]$ is the diffusion timestep, and $\bar{\alpha}^{(t)}$ is a hyperparameter controlling the signal/noise ratio. In practice, $\mathbf{Z}^{(t)}$ is sampled as $\mathbf{Z}^{(t)}=\sqrt{\bar{\alpha}^{(t)}}\mathbf{Z}^{(0)}+\sqrt{1-\bar{\alpha}^{(t)}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})$. Given the noised latents $\mathbf{Z}^{(t)}$, a diffusion model $\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},t)$ is trained to predict the added noise $\boldsymbol{\epsilon}$ by minimizing the MSE loss $\|\epsilon_{\theta}(\mathbf{Z}^{(t)},t)-\boldsymbol{\epsilon}\|^{2}$. Once trained, new latents $\hat{\mathbf{Z}}^{(0)}$ are sampled with the model $\boldsymbol{\epsilon}_{\theta}$ via iterative denoising [15], and the decoder $\mathcal{D}$ reconstructs the corresponding 3D molecule via $\hat{\mathbf{G}}=\mathcal{D}(\hat{\mathbf{Z}}^{(0)})$.
@@
Figure 3: Overview of the UDM-3D and UAE-3D models. The UAE-3D encodes 3D molecules from molecular space into a unified latent space, integrating multi-modal features such as atom types, chemical bonds, and 3D coordinates. Utilizing this latent space, UDM-3D employs a DiT to perform generative modeling. Then, the denoised latents are decoded back into 3D molecules.
@@ 3 Methodology
3.1 Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling
UAE-3D is designed to address the complexities of multi-modal and equivariance of molecular data by compressing the atomic features, bond features, and atomic coordinates into a unified latent space. This is achieved through three key components: (1) a Relatinoal-Transformer [21] that effectively integrates the multi-modal features into token sequences, (2) a Transformer-based [22] decoder and the reconstruction loss for 3D molecule reconstruction; and (3) our tailor-made SE(3)-equivariant data augmentations to train the model to learn 3D equivariance.
Compressing 3D Molecules with Relational Transformer (R-Trans) Encoder. Given a 3D molecular $\mathbf{G}=\langle\mathbf{F},\mathbf{E},\mathbf{X}\rangle$, we compute its initial atom-wise and edge-wise embeddings as:
$\mathbf{H}^{n}=\operatorname{MLP}([\mathbf{X};\mathbf{F}])\in\mathbb{R}^{|\mathcal{V}|\times d},$ (2) $\mathbf{H}^{e}=\operatorname{MLP}([\mathbf{E};\mathbf{D}])\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times d},$ (3)
where $\mathbf{D}_{ij}=\operatorname{GBF}(\|\mathbf{X}_{i}-\mathbf{X}_{j}\|^{2})\in\mathbb{R}^{d}$; $\operatorname{GBF}(\cdot)$ implements the Gaussian basis functions to expand inter-atomic distances as feature vectors [27]; and $d$ is the embedding dimension. The initial node embeddings $\mathbf{H}^{n}$ combines both atomic features and its 3D coordinates, while the edge embeddings $\mathbf{H}^{e}$ incorporate bond features and inter-atomic distances. $\mathbf{H}^{n}$ and $\mathbf{H}^{e}$ together effectively represent all 3D molecular information for the subsequent encoding. Given these embeddings, we process them with the molecular encoder $\mathcal{E}(\cdot)$ of $L$ layers of R-Trans [21]: